OpenStudy (jazzyfa30):

Use the angle sum or difference identity to find the exact value of each? cos 15 degrees A)- (squareroot6)-(squareroot2)/4 B)(squareroot6)+(squareroot2)/4 C)(squareroot6)-(squareroot2)/4 D)(squareroot2)-(squareroot6)/4

OpenStudy (jazzyfa30):

@ganeshie8

ganeshie8 (ganeshie8):

hint : 15 = 45-30

ganeshie8 (ganeshie8):

cos(15) = cos(45 - 30)

OpenStudy (jazzyfa30):

So i have to use my unit circle

ganeshie8 (ganeshie8):

apply cos(A - B) formula

ganeshie8 (ganeshie8):

yes u need to memorize sin/cos values for 45 and 30 degrees

OpenStudy (jazzyfa30):

cos45 = \[\sqrt{2}/2\] cos 30=\[\sqrt{3}/2\] sin45=\[\sqrt{2}/2\] sin 30= 1/2

ganeshie8 (ganeshie8):

excellent ! plug them in and simplify

OpenStudy (jazzyfa30):

isnt it subtraction

OpenStudy (jazzyfa30):

I got c

ganeshie8 (ganeshie8):

sorry i gave u the wrong formula :/

ganeshie8 (ganeshie8):

correct formula : cos(45 - 30) = cos45 cos30 \(\color{red}{+}\) sin45 sin30

ganeshie8 (ganeshie8):

please work it again lol, really sorry

OpenStudy (jazzyfa30):

ok so its the same just different symbol

ganeshie8 (ganeshie8):

yup !

OpenStudy (jazzyfa30):

can you help me with one more like that

ganeshie8 (ganeshie8):

sure :)